a machine, purchased for $10,000, has a depreciable life of 5 years. it will have an expected salvage value of $52,000 at the end of the depreciable life. using the double-declining balance method what is the book value at the end of year 3


1
Expert's answer
2021-04-30T10:48:07-0400

Depreciation rate is 100%/5years=20%. Using double-declining balance method, depreciation of the 1st year is 10,000×2×20%=4,000, the 2nd year - (10,000-4,000)×40%=2,400, 3rd year - (10,000-4,000-2,400)×40%=1,440. The book value at the end of 3rd year is 10,000-4,000-2,400-1,440=2,160.